Two particle of mass M and m moving in circular path having radius Rand r. If time period of two masses are same then what will be ratio of angular velocity

Grade:12

2 Answers

Amritpal Singh Saini
5 years ago
Time period is = 2pi/w , where w= angular velocity...... If the time periods are same , the we get w2=w1 OR the ratio of angular velocity is 1
Arun
5 years ago
The two particles m and M are moving in a circular motion with different radii  r and R..
Time period = T = 2 π / ω
=> T1/ T2 =  ω2 / ω1
=> ω2 / ω1 = 1  as  T1 = T2